OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through point (1, -2) and has a slope of 1/3 . 3x - y - 7 = 0 x - 3y + 7 = 0 x - 3y - 7 = 0

OpenStudy (sheraz12345):

c

OpenStudy (anonymous):

What is the equation of a line, in general form, that passes through points (-1, 2) and (5, 2)? y - 2 = 0 x - 2 = 0 y - x - 2 = 0

OpenStudy (cp9454):

y-2 = 0